# Cert-SSB: Toward Certified Sample-Specific Backdoor Defense

### v1

@@ III-A Preliminaries
The Main Pipeline of (Poisoning-based) Backdoor Attacks. Let $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ represents the benign dataset consisting of $n$ samples, where $\bm{x}_{i}\in\mathcal{X}$ is the $i$-th image, $y_{i}\in\mathcal{Y}=\{1,2,\cdots,K\}$ is its corresponding label, and $K$ denotes the total number of classes. In general, adversaries create a poisoned dataset $\mathcal{D}_{p}$ to train the target model using either a standard loss function or a customized one specified by the attacker. Specifically, $\mathcal{D}_{p}$ consists of two main parts: 1) the modified version of a selected subset ($i.e.$, $\mathcal{D}_{s}$) of $\mathcal{D}$, and 2) the remaining benign subset $\mathcal{D}_{b}$. Formally, $\mathcal{D}_{p}=\mathcal{D}_{m}(\mathcal{B}_{\bm{x}},\hat{y})\cup\mathcal{D}_{b}$, where $\mathcal{D}_{m}(\mathcal{B}_{\bm{x}},\hat{y})=\left\{\bm{x}_{i}+\bm{\Delta}_{i},\hat{y}\right\}_{i=1}^{d}$, $\mathcal{D}_{b}=\mathcal{D}\backslash\mathcal{D}_{s}=\left\{\bm{x}_{i},y_{i}\right\}_{i=d+1}^{n}$, $\mathcal{B}_{\bm{x}}$ introduces unique trigger patterns $\bm{\Delta}_{i}$ into the selected training instances, and $\hat{y}=G_{Y}(y)$. Here, $\lambda\triangleq\frac{|\mathcal{D}_{m}|}{|\mathcal{D}|}$ is the poisoning rate, and $G_{Y}$ is adversary-specified poisoned label generator. For example, in Badnets [32], $G_{Y}(y)=y_{t}$ for all-to-one attacks, where $y_{t}\in\mathcal{Y}$ is the target label, and $G_{Y}(y)=y+1$ mod $K$ for all-to-all attacks. The attack succeeds if the classifier predicts the target label $\hat{y}$ with high probability for a testing example $\bm{x}$ modified with the backdoor pattern $\mathcal{B}_{\bm{x}}$: $f(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}_{m}(\mathcal{B}_{\bm{x}},\hat{y}))=\hat{y}$.
Definition 1 (Boundary Samples and Closet Boundary Samples.).
@@
where $\alpha_{t}$ is a positive step size, $\bm{x}_{0}$ is an initial point satisfying $\phi_{y}(\bm{x}_{0};\bm{w})\leq 0$ and $\beta_{t}\in[0,1]$ is a line search parameter, to ensure $\bm{x}^{*}_{t+1}$ satisfies $\phi_{y}(\bm{x}^{*}_{t+1};\bm{w})=0$. In practice, $\bm{x}_{0}$ is randomly selected from the validation set, ensuring its label differs from $y$.
@@ III-B Analysis of Sample’s Distance to Decision Boundary
We hereby analyze how the distance from a sample to the decision boundary varies across inputs. Specifically, this distance is estimated using the ‘closest boundary sample’ defined in Definition 1 to avoid the inaccurate estimation using a random boundary one since there are multiple of them.
Setting. We hereby use BadNets [32] attack with a ResNet model [63] on the CIFAR-10 [64] datasets for discussion. Specifically, we set the target label $y_{t}$ as ‘0’ and the poisoning rate as 5%. Following the previous work [24], we use a one-pixel patch located at the lower right corner of the image as the trigger pattern. We randomly select 2,000 poisoned samples and use Eq. (6) to generate their closest boundary samples for the target label $y_{t}$. During training, we compute the $\ell_{2}$ norm between each sample and its closest decision boundary sample. Samples with a small distance to the decision boundary are referred to as easy samples, while those with a larger distance are referred to as hard samples.
Result. As shown in Figure 2, the $\ell_{2}$ distances to the closest boundary samples vary significantly among different samples in the poisoned dataset. Specifically, although most samples have relatively small distances ($e.g.$, $\ell_{2}\leq 0.3$), a considerable number of hard samples exhibit larger distances to their closest boundary samples. Therefore, these hard samples require a larger magnitude of noise to effectively defend against backdoor attacks during training. In contrast, for easy samples with smaller distances, only a smaller magnitude of noise is needed to achieve the desired defense effect. For samples that fall between hard samples and easy samples, a trade-off and adjustment in noise selection are necessary.
@@ III-C Limitations of Fixed Noise in Testing Samples
Result. As shown in Figures 3(a) and 3(b), the three samples from both MNIST and CIFAR-10 datasets exhibit a trend where the certification radius first increases and then decreases as the noise magnitude increases. Notably, although the model was trained with $\sigma=1.0$, the optimal certification radius does not occur at this noise level. Taking the MNIST dataset as an example, sample 1 reaches its maximum certification radius of approximately 1.5 at $\sigma=0.8$, sample 2 peaks at about 1.3 when $\sigma$ approaches 0.9, while sample 3 maintains a relatively stable low value. This result suggests that the optimal certification performance is not necessarily achieved by using the same noise magnitude during testing as in training. Therefore, the $\sigma$ value should be optimized for each sample to achieve the maximum certification radius.
@@ III-D Limitations of Fixed Noise in Training Samples
Figure 4: The main pipeline of our Cert-SSB consists of two stages. In the first stage, we adopt a stochastic gradient ascent (SGA) strategy to iteratively optimize the noise in order to maximize the certification radius $r$, thereby solving for the optimal noise ($i.e.$, sample-specific noise). The value of $r$ is computed based on the predictions of a base model trained with fixed noise. This optimized noise is then injected into the poisoned training set and used to train $M$ smoothed models. In the second stage, the $M$ smoothed models trained in the first stage are aggregated to generate the final smoothed prediction. Notably, under this setting, the traditional certification method, which typically assumes a fixed noise level, is no longer applicable. To conquer this challenge, we propose a novel storage-update-based certification method, which ensures that each certification region is non-overlapping and maintains consistent predictions within each region (see Figure 5 for more details).
@@
Result. As shown in Figures 3(c) and 3(d), the certification radii of different samples exhibit distinct trends as the noise level varies. Overall, while some samples achieve larger certification radii at appropriate noise levels, others are more sensitive to noise, showing instability or even misclassification at higher noise values. For example, Sample 1 shows a continuously increasing certification radius as the noise $\sigma$ increases, indicating that its robustness remains stable even at higher noise levels. Sample 2 exhibits a stable certification radius in the range of $\sigma=0.6$ to $0.8$, without significant changes as the noise level further increases. This suggests that this noise level may be optimal for this sample. In this case, increasing the noise further may not improve the certification radius and could even negatively impact classification accuracy. Therefore, for this sample, a trade-off must be made between accuracy and robustness. In contrast, Sample 3 experiences a gradual decrease in certification radius as the noise level increases and eventually undergoes misclassification at higher noise levels. The “$\times$” markers in the figure indicate misclassified points. This result suggests that the noise level used for training should be optimized based on the characteristics of individual samples rather than using a fixed value to achieve better certification performance.
@@ IV-A2 Goal of Certified Defense
The primary goal is to defend against poison-only backdoor attacks by obtaining a robustness threshold $r$ through analysis, ensuring that if the total backdoor modification satisfies $\sqrt{\sum_{i=1}^{d}\left\|\bm{\Delta}_{i}\right\|_{2}^{2}}<r$, the classifier’s predictions on testing samples containing backdoor triggers remain unaffected by whether the model was trained on poisoned or clean data. In other words, the model’s predictions should be consistent, expressed as: $f(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}_{m}(\mathcal{B}_{\bm{x}},\hat{y}))=f(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}_{m}(\emptyset))$, where $\mathcal{D}_{m}(\emptyset)$ denotes the clean dataset without any embedded backdoor patterns ($i.e.$, $\bm{\Delta}_{i}=0$).
@@ IV-B Overview of the Proposed Method
As demonstrated in Section III, existing randomized smoothing-based certified backdoor defenses exhibit suboptimal certification performance, regardless of whether fixed noise is applied to training or testing samples. This is because each sample has a different distance to the decision boundary. To address this issue, we propose a sample-specific certified backdoor defense method, dubbed Cert-SSB, in which the noise level is adaptively adjusted for each individual sample.
As shown in Figure 4, our method consists of two main stages: (1) Cert-SSB training stage, and (2) Cert-SSB inference stage. In the training stage, we apply stochastic gradient ascent to iteratively solve for the optimal noise level $\sigma_{\bm{x}}^{*}$ that maximizes the certification radius. The radius is computed based on the prediction probabilities of a base classifier trained with fixed noise. Once obtained, the sample-specific noise $\sigma_{\bm{x}}^{*}$ is injected into the poisoned training set to train $M$ smoothed models. In the inference stage, we aggregate predictions from the $M$ smoothed models trained in the first stage to produce the final smoothed output. Intuitively, as long as the predicted probability of the most likely class exceeds that of the second most likely class, the model can be considered certifiably robust. However, under this sample-specific noise setting, traditional certification methods become inapplicable, as they typically assume a uniform noise level across all inputs. To overcome this limitation, we introduce a storage-update-based certification method, which categorizes certification regions ($i.e.$, regions defined by the certified radius of each input) to ensure that they remain non-overlapping across inputs and that prediction consistency is maintained within each region (see Figure 5). The technical details are as follows.
IV-C Cert-SSB Training: Train the Model with Optimized Noise
In this stage, we train multiple smoothed models based on optimized noise, where the optimized noise is obtained by applying stochastic gradient ascent to maximize the certification radius. In general, the certification radius is computed based on the predictions of a classifier trained with fixed noise.
@@ IV-C1 Optimized Sample-Specific Noise Generation
Given a smoothed classifier with fixed noise $\sigma_{0}$ (which reduces to the base classifier $f(\bm{x},\mathcal{D})$ when $\sigma_{0}=0$), our goal is to construct a new smoothed classifier $g(\bm{x},\mathcal{D},\sigma_{\bm{x}}^{*})$ based on an optimized noise $\sigma_{\bm{x}}^{*}$ for each sample. The new classifier should ensure that for all samples $\bm{x}$, the predictions of the two smoothed classifiers (with $\sigma_{0}$ and $\sigma_{\bm{x}}^{*}$) remain identical while also maximizing the certification radius at each sample $\bm{x}$. Formally, let $y_{A}$ represent the predicted label under the fixed noise level $\sigma_{0}$, it can be defined as:
@@
where $Z\sim\mathcal{N}(0,\sigma_{0}^{2}I)$ is assumed to be independent, and $D\sim\mathcal{N}(0,\sigma_{0}^{2}I)$ consists of $n$ i.i.d. random variables $D^{(i)}$, each added to a training instance in $\mathcal{D}$. For each input $\bm{x}$, the optimized noise $\sigma_{\bm{x}}^{*}$ is obtained by solving for $\sigma$ that maximizes the certification radius $r(\bm{x},\sigma)$ in Eq. (4):
@@
where $P_{A}(\bm{x},\mathcal{D},\sigma)=\mathcal{P}_{\bm{\epsilon}(Z,D)}[f^{y_{A}}(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)]$, $P_{B}(\bm{x},\mathcal{D},\sigma)=\max_{y\neq y_{A}}\mathcal{P}_{\bm{\epsilon}(Z,D)}[f^{y}(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)]$.
@@
$\begin{array}[]{ll}&\bigtriangledown_{\sigma^{t}}\{\frac{\sigma^{t}}{2}\cdot[\Phi^{-1}(\frac{1}{J}{\textstyle\sum_{j=1}^{J}f^{y_{A}}(\bm{x}+\mathcal{B}_{\bm{x}}+Z_{i},\mathcal{D}+\bm{\delta}+D_{i})})\\ &-\Phi^{-1}(\max_{y\neq y_{A}}\frac{1}{J}{\textstyle\sum_{j=1}^{J}f^{y}(\bm{x}+\mathcal{B}_{\bm{x}}+Z_{i},\mathcal{D}+\bm{\delta}+D_{i})})]\},\end{array}$ (9)
@@
where $\hat{P_{A}}(\bm{x},\mathcal{D},\sigma)=\mathcal{P}_{\bm{\epsilon}(\hat{Z},\hat{D})}[f^{y_{A}}(\bm{x}+\mathcal{B}_{\bm{x}}+\sigma\hat{Z},\mathcal{D}+\bm{\delta}+\sigma\hat{D})]$, $\hat{P_{B}}(\bm{x},\mathcal{D},\sigma)=\max_{y\neq y_{A}}\mathcal{P}_{\bm{\epsilon}(\hat{Z},\hat{D})}[f^{y}(\bm{x}+\mathcal{B}_{\bm{x}}+\sigma\hat{Z},\mathcal{D}+\bm{\delta}+\sigma\hat{D})]$. Note that under this reparameterization, the distributions $\hat{Z}$ and $\hat{D}$ are no longer dependent on the optimization variable $\sigma$. As a result, Eq. (10) typically yields lower-variance gradient estimates compared to the original formulation in Eq. (8).

### v2

@@ III-A Preliminaries
The Main Pipeline of (Poisoning-based) Backdoor Attacks. Let $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ represents the benign dataset consisting of $n$ samples, where $\bm{x}_{i}\in\mathcal{X}$ is the $i$-th image, $y_{i}\in\mathcal{Y}=\{1,2,\cdots,K\}$ is its corresponding label, and $K$ denotes the total number of classes. In general, adversaries create a poisoned dataset $\mathcal{D}_{p}$ to train the target model using either a standard loss function or a customized one specified by the attacker. Specifically, $\mathcal{D}_{p}$ consists of two main parts: 1) the modified version of a selected subset ($i.e.$, $\mathcal{D}_{s}$) of $\mathcal{D}$, and 2) the remaining benign subset $\mathcal{D}_{b}$. Formally, $\mathcal{D}_{p}=\mathcal{D}_{m}(\bm{\delta},\hat{y})\cup\mathcal{D}_{b}$, where $\mathcal{D}_{m}(\bm{\delta},\hat{y})=\left\{\bm{x}_{i}+\bm{\Delta}_{i},\hat{y}\right\}_{i=1}^{\tilde{r}}$, $\mathcal{D}_{b}=\mathcal{D}\backslash\mathcal{D}_{s}=\left\{\bm{x}_{i},y_{i}\right\}_{i=\tilde{r}+1}^{n}$, $\bm{\delta}$ denotes the collection of unique trigger patterns $\bm{\Delta}_{i}$ injected into the selected training instances, and $\hat{y}=G_{Y}(y)$. Here, $\lambda\triangleq\frac{|\mathcal{D}_{m}|}{|\mathcal{D}|}$ is the poisoning rate, and $G_{Y}$ is adversary-specified poisoned label generator. For example, in Badnets [19], $G_{Y}(y)=y_{t}$ for all-to-one attacks, where $y_{t}\in\mathcal{Y}$ is the target label, and $G_{Y}(y)=y+1$ mod $K$ for all-to-all attacks. The attack succeeds if the classifier predicts the target label $\hat{y}$ for a testing example $\bm{x}$ modified with the backdoor pattern $\mathcal{B}_{\bm{x}}$: $f(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}_{p})=\hat{y}$.
Definition 1 (Boundary Samples and Closest Boundary Samples).
@@
where $\alpha_{t}$ is a positive step size, $\bm{x}_{0}$ is an initial point s.t. $\phi_{y}(\bm{x}_{0};\bm{w})\leq 0$ and $\beta_{t}\in[0,1]$ is a line search parameter s.t. $\phi_{y}(\bm{x}^{*}_{t+1};\bm{w})=0$. In practice, $\bm{x}_{0}$ is randomly selected from the validation set, ensuring its label differs from $y$.
@@ III-B Analysis of Sample’s Distance to Decision Boundary
We hereby analyze how the distance from a poisoned sample to the decision boundary of the target class varies across different inputs. Here, the closest decision boundary refers to the minimal perturbation required to change the prediction from the target label $y_{t}$ to any non-target class $y\neq y_{t}$. Specifically, this distance is estimated using the ‘closest boundary sample’ defined in Definition 1 to avoid the inaccurate estimation using a random boundary one since there are multiple of them.
Setting. We hereby use BadNets [19] attack with a ResNet model [69] on the CIFAR-10 [30] datasets for discussion. Specifically, we set the target label $y_{t}$ as ‘0’ and the poisoning rate as 5%. Following the previous work [59], we use a one-pixel patch located at the lower right corner of the image as the trigger pattern. We randomly select 2,000 poisoned testing samples ($i.e.$, samples containing the trigger and predicted as the target label $y_{t}$ by the backdoored model) and use Eq. (6) to generate their closest boundary samples for the target label $y_{t}$. We then compute the $\ell_{2}$ norm between each poisoned sample and its closest boundary sample. Samples with a small distance are referred to as easy poisoned samples, while those with a larger distance are referred to as hard poisoned samples.
Figure 3: Distribution of $\ell_{2}$ norm distances between poisoned samples and their closest boundary samples. Samples closer to the boundary are ‘easy’ poisoned samples, while those farther away are ‘hard’ poisoned samples.
Result. As shown in Figure 3, the $\ell_{2}$ distances to the closest boundary samples vary significantly among different samples in the poisoned dataset. Specifically, although most samples have relatively small distances ($e.g.$, $\ell_{2}\leq 0.3$), a considerable number of hard samples exhibit larger distances to their closest boundary samples. Therefore, these hard poisoned samples would require a larger magnitude of noise to effectively suppress the backdoor effect. In contrast, for easy poisoned samples with smaller distances, only a smaller magnitude of noise is needed to achieve the desired defense effect. For samples that fall between easy and hard poisoned samples, a trade-off and adjustment in noise selection are necessary.
@@ III-C Limitations of Fixed Noise in Testing Samples
Result. As shown in Figures 2(a) and 2(b), the three samples from both MNIST and CIFAR-10 datasets exhibit a trend where the certification radius first increases and then decreases as the noise magnitude increases. Notably, although the model was trained with $\sigma=1.0$, the optimal certification radius does not occur at this noise level. Taking the MNIST dataset as an example, sample 1 reaches its maximum certification radius of approximately 1.5 at $\sigma=0.8$, sample 2 peaks at about 1.3 when $\sigma$ approaches 0.9, while sample 3 maintains a relatively stable low value. This result suggests that the optimal certification performance is not necessarily achieved by using the same noise magnitude during testing as in training. Therefore, the $\sigma$ value should be optimized for each sample to achieve the maximum certification radius.
@@ III-D Limitations of Fixed Noise in Training Samples
Figure 4: The main pipeline of our Cert-SSBD consists of two stages. In the first stage, we adopt a stochastic gradient ascent (SGA) to iteratively optimize the noise to maximize the certification radius $r$, thereby obtaining optimal sample-specific noise to train $M$ smoothed models. In the second stage, the $M$ smoothed models trained in the first stage are aggregated to generate the final prediction. In particular, we propose a storage-update-based certification method to ensure non-overlapping certification regions and consistent predictions under the sample-specific noise setting (see Figure 5 for more details).
@@
Result. As shown in Figures 2(c) and 2(d), the certification radii of different samples exhibit distinct trends as the noise level varies. Overall, while some samples achieve larger certification radii at appropriate noise levels, others are more sensitive to noise, showing instability or even misclassification at higher noise values. For example, Sample 1 shows a continuously increasing certification radius as the noise $\sigma$ increases, indicating that its robustness remains stable even at higher noise levels. Sample 2 exhibits a stable certification radius in the range of $\sigma=0.6$ to $0.8$, without significant changes as the noise level further increases. This suggests that this noise level may be optimal for this sample. In this case, increasing the noise further may not improve the certification radius and could even negatively impact classification accuracy. Therefore, for this sample, a trade-off must be made between accuracy and robustness. In contrast, Sample 3 experiences a gradual decrease in certification radius as the noise level increases and eventually undergoes misclassification at higher noise levels. The $\times$” markers in the figure indicate misclassified points. This result suggests that the noise level used for training should be optimized based on the characteristics of individual samples rather than using a fixed value to achieve better certification performance.
@@ IV-A2 Goal of Certified Defense
The primary goal is to defend against poison-only backdoor attacks by obtaining a robustness threshold $r$ through analysis, ensuring that if the total backdoor modification satisfies $\sqrt{\sum_{i=1}^{n}\left\|\bm{\Delta}_{i}\right\|_{2}^{2}}<r$, the classifier’s predictions on testing samples containing backdoor triggers remain unaffected by whether the model was trained on poisoned or clean data. In other words, the model’s predictions should be consistent, expressed as: $f(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}_{p})=f(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D})$, where $\mathcal{D}_{p}=\mathcal{D}_{m}(\bm{\delta},\hat{y})\cup\mathcal{D}_{b}$ is the poisoned training set and $\mathcal{D}$ is the clean training set ($i.e.$, $\bm{\Delta}_{i}=\mathbf{0}$ for all $i$).
@@ IV-B Overview of the Proposed Method
As demonstrated in Section III, existing randomized smoothing-based certified backdoor defenses exhibit suboptimal certification performance, regardless of whether fixed noise is applied to training or testing samples. This is because each sample has a different distance to the decision boundary. To address this issue, we propose a sample-specific certified backdoor defense method, dubbed Cert-SSBD, in which the noise level is adaptively adjusted for each individual sample.
As shown in Figure 4, our method consists of two main stages: (1) Cert-SSBD training stage, and (2) Cert-SSBD inference stage. In the training stage, we apply stochastic gradient ascent to iteratively solve for the optimal noise level $\sigma_{\bm{x}}^{*}$ that maximizes the certification radius. Once the sample-specific noise $\sigma_{\bm{x}}^{*}$ is obtained, it is injected into the poisoned training set to train $M$ smoothed models. In the inference stage, we aggregate the predictions of these $M$ smoothed models to generate the final output. Intuitively, as long as there exists a non-trivial gap between the predicted probabilities of the most likely class and the runner-up class, a non-zero certification radius can be obtained, and the model can be considered certifiably robust. However, under this sample-specific noise setting, traditional certification methods become inapplicable, as they typically assume a uniform noise level across all inputs. To overcome this limitation, we introduce a storage-update-based certification method. This method categorizes certification regions ($i.e.$, regions defined by the certified radius of each input) to ensure that these regions remain non-overlapping across different inputs and maintain prediction consistency within each region (see Figure 5 for details). The technical details are as follows.
IV-C Cert-SSBD Training: Train Models with Optimized Noises
In this stage, we describe the training stage of Cert-SSBD, which consists of two sequential steps: 1) optimizing a sample-specific noise scale $\sigma_{\bm{x}}^{*}$ via stochastic gradient ascent (SGA) to maximize the certified radius, and 2) training an ensemble of smoothed models using the resulting optimized noise scales. In general, the certified radius is computed from the class probabilities of a smoothed classifier trained with a fixed initialized noise level.
@@ IV-C1 Optimized Sample-Specific Noise Generation
Given a base smoothed classifier with a fixed noise level $\sigma_{0}$ ($i.e.$, a predefined initialized noise scale), our goal is to construct a new smoothed classifier $g(\bm{x},\mathcal{D}_{p},\sigma_{\bm{x}}^{*})$ based on a set of optimized, sample-specific noise scales $\{\sigma_{\bm{x}_{i}}^{*}\}_{i=1}^{n}$, where each $\sigma_{\bm{x}_{i}}^{*}$ corresponds to a single training sample. Note that when $\sigma_{0}=0$, the base smoothed classifier degenerates to the base classifier $f(\bm{x},\mathcal{D}_{p})$. The new classifier should ensure that for all training samples $\bm{x}$, the predictions of the two smoothed classifiers (with $\sigma_{0}$ and $\sigma_{\bm{x}}^{*}$) remain identical, while also maximizing the certification radius for each sample. Formally, we define $y_{A}$ as the most probable (top-1) predicted class under the fixed noise level $\sigma_{0}$, and $y_{B}$ as the runner-up class with the second-highest predicted probability, $i.e.$,
@@
where $Z\sim\mathcal{N}(0,\sigma_{0}^{2}I)$ is assumed to be independent, and $D\sim\mathcal{N}(0,\sigma_{0}^{2}I)$ consists of $n$ i.i.d. random variables $D^{(i)}$, each added to a training instance in $\mathcal{D}$. The optimized noise $\sigma_{\bm{x}}^{*}$ is obtained by solving a sample-specific optimization problem that maximizes the certification radius $r(\bm{x},\sigma)$ in Eq. (4):
@@
where $P_{A}(\bm{x},\mathcal{D},\sigma)=\mathcal{P}_{\bm{\epsilon}(Z,D)}[f^{y_{A}}(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)]$, $P_{B}(\bm{x},\mathcal{D},\sigma)=\max_{y_{B}\neq y_{A}}\mathcal{P}_{\bm{\epsilon}(Z,D)}[f^{y}(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)]$.
@@
$\begin{array}[]{ll}&\bigtriangledown_{\sigma^{t}}\{\frac{\sigma^{t}}{2}\cdot[\Phi^{-1}(\frac{1}{J}{\textstyle\sum_{j=1}^{J}f^{y_{A}}(\bm{x}+\mathcal{B}_{\bm{x}}+Z_{i},\mathcal{D}+\bm{\delta}+D_{i})})\\ &-\Phi^{-1}(\max_{y_{B}\neq y_{A}}\frac{1}{J}{\textstyle\sum_{j=1}^{J}f^{y}(\bm{x}+\mathcal{B}_{\bm{x}}+Z_{i},\mathcal{D}+\bm{\delta}+D_{i})})]\},\end{array}$ (9)
@@
[…]